OpenStudy (anonymous):

convert 1.50μm2 to square meters

OpenStudy (anonymous):

\[1 \mu m=10^{-6}m \rightarrow 1\mu m^2=10^{-12}m^2\]\[1.5 \mu m^2=1.5·10^{-12}m^2\]